OpenStudy (anonymous):

Which of the following has a graph that is a straight line? (1 to 5 Questions)

OpenStudy (anonymous):

(Question1) Which of the following has a graph that is a straight line? Equation 1: y = 2x + 7 Equation 2: y2 = x − 1 Equation 3: y = 2x2 + 4 Equation 4: y = 3x3